The weights, in pounds, of eight students in a class are:

128
193
166
147
202
183
181
158

Using the data above, what is the standard error of the sample mean? Answer choices are rounded to the hundredths place.
Answer #1

AN From above data. n=8 = 169.15. set ( *: - 1)* = 24-77 Standard 24.77 espon ewoon = = 2477 - 8.76 200